# Get rich quick: exact solutions reveal how unbalanced initializations promote rapid feature learning

### v1

@@ C.1 Surface of a piecewise linear network
For a piecewise linear network $f(x;\theta)$, comprising $m$ hidden neurons with pre-activation $z_{i}(x;\theta)$ for $i\in[m]$, let the activation pattern $\mathcal{A}$ represent an assignment of signs $a_{i}\in\{-1,1\}$ to each hidden neuron. The activation region $\mathcal{R}(\mathcal{A};\theta)$ is the subset of input space that generates $\mathcal{A}$,
@@
The input-output map is linear within each non-empty activation region and continuous at the boundary between regions. Linearity implies that every non-empty [^9] While it is trivial to see that for a network $f(x;\theta)$ with $m$ hidden neurons there are $2^{m}$ distinct activation patterns, not all activation patterns are attainable. See Raghu et al. [63] for a discussion. activation region is associated with a linear predictor vector $\beta_{\mathcal{R}}\in\mathbb{R}^{d}$ such that for all $x\in R(\mathcal{A};\theta)$, $\beta_{\mathcal{R}}=\nabla_{x}f(x;\theta)$. Continuity implies that the boundary between regions is formed by a hyperplane determined by where the pre-activation for a neuron is exactly zero, $\{x:z_{i}(x;\theta)=0\}$. When the neighboring regions have different linear predictors [^10] It is possible for neighboring regions to have the same linear predictor. Some works define linear regions as maximally connected component of input space with the same linear predictor [81]., then this hyperplane is orthogonal to their difference, which is a vector in the span of the first-layer weights. Taken together, this implies that the union of all activation regions forms a convex partition of input space, as shown in Fig. 9. We now present a surprisingly simple, yet to the best of our knowledge not previously understood property of this partition:
@@
The proof of this proposition is straightforward. There is one color for regions with an even number of active neurons and another for regions with an odd number of active neurons. Because $f(x;\theta)$ lacks redundant neurons, there does not exist a boundary between activation regions where two neurons activations change simultaneously. In this work, we solely utilize this proposition for visualization purposes, as shown in Fig. 9. Nonetheless, we believe it may be of independent interest as it strengthens the connection between the surface of piecewise linear networks and the mathematics of paper folding, a connection previously alluded to in the literature [76].
@@ C.2 Dynamics of a two-layer piecewise linear network
We consider the dynamics of a two-layer piecewise linear network without biases, $f(x;\theta)=a^{\intercal}\sigma(Wx)$, where $W\in\mathbb{R}^{h\times d}$ and $a\in\mathbb{R}^{h}$. The activation function is $\sigma(z)=\max\{z,\gamma z\}$ for $\gamma\in[0,1)$, which includes ReLU $\gamma=0$ and Leaky ReLU $\gamma\in(0,1)$. We permit $h>d$, which in the limit as $h\to\infty$, ensures the network possesses the functional expressivity to represent any continuous nonlinear function from $\mathbb{R}^{d}$ to $\mathbb{R}$ passing through the origin. Following a similar strategy used in Section 4, we consider the contribution to the input-output map from a single hidden neuron $i\in[h]$ with parameters $w_{i}\in\mathbb{R}^{d}$ and $a_{i}\in\mathbb{R}$. As in the linear setting, each hidden neuron is associated with a conserved quantity, $\delta_{i}=\eta_{w}a_{i}^{2}-\eta_{a}\|w_{i}\|^{2}$. Unlike in the linear setting, this neuron’s contribution to the output $f(x_{j};\theta)$ is regulated by whether the input $x_{j}$ is in the neuron’s active halfspace, $\{x\in\mathbb{R}^{d}:w_{i}^{\intercal}x>0\}$. Let $C\in\mathbb{R}^{h\times n}$ be the matrix with elements $c_{ij}=\sigma^{\prime}(w_{i}^{\intercal}x_{j})$, which determines the activation of the $i^{\mathrm{th}}$ neuron for the $j^{\mathrm{th}}$ training data point. The subgradient $\sigma^{\prime}(z)=1$ if $z>0$, $\sigma^{\prime}(z)\in[\gamma,1]$ if $z=0$, and $\sigma^{\prime}(z)=\gamma$ if $z<0$. These activation functions exhibit positive homogeneity, implying $\sigma(z)=\sigma^{\prime}(z)z$. Thus, we can express $\sigma(w_{i}^{\intercal}x_{j})=c_{ij}w_{i}^{\intercal}x_{j}$, allowing us to express the gradient flow dynamics for $w_{i}$ and $a_{i}$ as
$\dot{a}_{i}=-\eta_{a}w_{i}^{\intercal}\left(\sum_{j=1}^{n}c_{ij}x_{j}\rho_{j}\right),\qquad\dot{w}_{i}=-\eta_{w}a_{i}\left(\sum_{j=1}^{n}c_{ij}x_{j}\rho_{j}\right),$ (124)
@@
where $\rho_{j}=f(x_{j};\theta)-y_{j}$ is the residual associated with the $j^{\mathrm{th}}$ training data point. If we let $\beta_{i}=a_{i}w_{i}$, which determines the contribution of each hidden neuron to the output $f(x_{j};\theta)$, then its not hard to see that the gradient flow dynamics of $\beta_{i}$ are
$\dot{\beta}_{i}=-\underbrace{\left(\eta_{w}a_{i}^{2}I_{d}+\eta_{a}w_{i}w_{i}^{\intercal}\right)}_{M_{i}}\underbrace{\left(\textstyle\sum_{j=1}^{n}c_{ij}x_{j}\rho_{j}\right)}_{\xi_{i}}.$ (125)
@@
As in the linear setting, the matrix $M_{i}\in\mathbb{R}^{d\times d}$ appears as a preconditioning matrix on the dynamics Using the exact same derivation presented in Section A.3, whenever $a_{i}^{2}>0$, we can express $M_{i}$ entirely in terms of $\beta_{i}$ and $\delta_{i}$,
$M_{i}=\frac{\sqrt{\delta_{i}^{2}+4\eta_{a}\eta_{w}\|\beta_{i}\|^{2}}+\delta_{i}}{2}I_{d}+\frac{\sqrt{\delta_{i}^{2}+4\eta_{a}\eta_{w}\|\beta_{i}\|^{2}}-\delta_{i}}{2}\frac{\beta_{i}\beta_{i}^{\intercal}}{\|\beta_{i}\|^{2}}.$ (126)
@@
However, unlike in the linear setting, the vector $\xi_{i}\in\mathbb{R}^{d}$ driving the dynamics is not shared for all neurons because of its dependence on $c_{ij}$. Additionally, the NTK matrix in this setting depends on $M_{i}$ and $C$, with elements $K_{jk}=\sum_{i=1}^{h}c_{ij}x_{j}^{\intercal}\left(\eta_{w}a_{i}^{2}I_{d}+\eta_{a}w_{i}w_{i}^{\intercal}\right)x_{k}c_{ik}$. Thus, in order to asses the temporal dynamics of the NTK matrix, we must understand the dynamics of $M_{i}$ and $C$. We consider a signed spherical coordinate transformation separating the dynamics of $\beta_{i}$ into its directional $\hat{\beta}_{i}=\mathrm{sgn}(a_{i})\tfrac{\beta_{i}}{\|\beta_{i}\|}$ and radial $\mu_{i}=\mathrm{sgn}(a_{i})\|\beta_{i}\|$ components, such that $\beta_{i}=\mu_{i}\hat{\beta}_{i}$. Here, $\hat{\beta}_{i}$ determines the orientation and direction of the halfspace where the $i^{\mathrm{th}}$ neuron is active, while $\mu_{i}$ determines the slope of the linear region in this halfspace. These coordinates evolve according to,
@@
$\dot{\mu}_{i}=-\sqrt{\delta_{i}^{2}+4\eta_{a}\eta_{w}\mu_{i}^{2}}\hat{\beta}_{i}^{\intercal}\xi_{i},\qquad\dot{\hat{\beta}}_{i}=-\frac{\sqrt{\delta_{i}^{2}+4\eta_{a}\eta_{w}\mu_{i}^{2}}+\delta_{i}}{2\mu_{i}}\left(I_{d}-\hat{\beta}_{i}\hat{\beta}_{i}^{\intercal}\right)\xi_{i}.$ (127)
@@
These equations can be derived directly from Eq. 124 through chain rule similar to Section A.2.1. In fact its worth noting that the this change of coordinates is very similar to the change of coordinates used in the single-neuron analysis. Expressed in terms of the parameters, $\hat{\beta}_{i}=\frac{w_{i}}{\|w_{i}\|}$ and $\mu_{i}=a_{i}\|w_{i}\|$.
@@ D.1 Figure 1: Teacher-Student with Two-layer ReLU Networks
Note that the base student initialization described thus far is perfectly balanced at each neuron, that is $\delta_{i}=0$ for $i\in[m]$; we also define this to be our setting where the scale $\tau$ is 1. In order to transform the base initialization into a particular setting of $\tau$ and $\delta$, we first solve for the relative layer scaling $\alpha$ in $\delta^{2}=\tau^{2}(\alpha^{2}-\alpha^{-2})$ and then scale each $w_{i}$ by $\tau/\alpha$ and each $a_{i}$ by $\tau\alpha$.
[…]
@@
Here the setting is: $d=100$, $m=50$, $k=3$, and $n=1000$, as in Fig. 1c of [8]. Training is performed with learning rate of 5e-3$/\tau^{2}$. Test error is computed as mean square error over a held-out set of 10,000 datapoints.
We sweep over $\tau$ over a logarithmic scale in the range $[0.1,2]$ and $\delta$ over a linear scale in the range $[-1,1]$. We average over 16 random seeds, where the seed controls the sampling of: the teacher weights $\theta^{\mathrm{teacher}}$, the base initialization of $\theta^{\mathrm{student}}$, and the training data $\{x^{(i)}\}_{i=1}^{n}$. In this way, each random seed is used for a sweep over all combinations of $\tau$ and $\delta$ in the sweep; we simply apply the scaling described above to get to each point on the $(\tau,\delta)$ grid.
The kernel distance computed is as defined in [25], where here we compute it at time $t$ relative to the kernel at initialization, i.e. $S(t)=1-\langle K_{0},K_{t}\rangle/\left(\|K_{0}\|_{F}\|K_{t}\|_{F}\right)$. In Fig. 10, we additionally plot Hamming and parameter distances relative to initialization, as well as training loss, while training for ten times longer than in Fig. 1 (b).
@@ D.2 Figures 2, 3, 4: Single-Neuron Linear Network
Figures 2, 3, and 4 were generated by simulating gradient flow using scipy.integrate.solve_ivp function with the RK45 method for solving the ODEs, with a relative tolerance of $1\times 10^{-6}$ and time span of $(0,20)$. In the experiments with full-rank data, we used $X^{\intercal}X=\mathbf{I}_{2}$, $\beta_{*}=\left[\begin{smallmatrix}0\\ 1\end{smallmatrix}\right]$, and $\beta_{0}=\left[\begin{smallmatrix}-1\\ 0\end{smallmatrix}\right]$. For the experiment with low-rank data, we used $X^{\intercal}X=\left[\begin{smallmatrix}0.25&0.5\\ 0.5&1\end{smallmatrix}\right]$, $\beta_{*}=\left[\begin{smallmatrix}0.44\\ 0.88\end{smallmatrix}\right]$, and $\beta_{0}=\left[\begin{smallmatrix}0.4\\ 0.05\end{smallmatrix}\right]$. See the discussion in Section A.2 for details on how we determined our theoretical predictions. A notebook generating all the figures is provided.
@@ D.3 Figure 5:
We are training a small ResNet based on the CIFAR10 script provided in the DAWN benchmark. [^11] Code available here. The only modification to the provided code base is that we set the weight decay parameter to 00, since this might confound our results. Moreover, we are dividing the convolutional filters weights by a parameter $\alpha$ (after standard initialization) which controls the balancedness of the network. To quantify the smoothness of the filters, we compute the normalized Laplacian of each filter $w_{ij}\in\mathbb{R}^{15\times 15}$, over input $i=(1,2,3)$ and output $j=(1,...,64)$ channels
@@
We are training a one layer transformer model on the modular arithmetic task in Power et al. [66]. Our experimental code is based on an existing Pytorch implementation. [^12] Code available here. The only modifications to the provided code base is that we use a single transformer layer (instead of the default 2-layer model). Prior analysis in Nanda et al. [69] has shown that this model can learn a minimal (attention-based) circuit that solves the task.

### v2

@@ C.1 Surface of a piecewise linear network
For a piecewise linear network $f(x;\theta)$, comprising $N$ hidden neurons with pre-activation $z_{i}(x;\theta)$ for $i\in[N]$, let the activation pattern $\mathcal{A}$ represent an assignment of signs $a_{i}\in\{-1,1\}$ to each hidden neuron. The activation region $\mathcal{R}(\mathcal{A};\theta)$ is the subset of input space that generates $\mathcal{A}$,
@@
The input-output map is linear within each non-empty activation region and continuous at the boundary between regions. Linearity implies that every non-empty [^9] While it is trivial to see that for a network $f(x;\theta)$ with $N$ hidden neurons there are $2^{N}$ distinct activation patterns, not all activation patterns are attainable. See Raghu et al. [65] for a discussion. activation region is associated with a linear predictor vector $\beta_{\mathcal{R}}\in\mathbb{R}^{d}$ such that for all $x\in R(\mathcal{A};\theta)$, $\beta_{\mathcal{R}}=\nabla_{x}f(x;\theta)$. Continuity implies that the boundary between regions is formed by a hyperplane determined by where the pre-activation for a neuron is exactly zero, $\{x:z_{i}(x;\theta)=0\}$. When the neighboring regions have different linear predictors [^10] It is possible for neighboring regions to have the same linear predictor. Some works define linear regions as maximally connected component of input space with the same linear predictor [87]., then this hyperplane is orthogonal to their difference, which is a vector in the span of the first-layer weights. Taken together, this implies that the union of all activation regions forms a convex partition of input space, as shown in Fig. 9. We now present a surprisingly simple, yet to the best of our knowledge not previously understood property of this partition:
@@
The justification for this proposition is straightforward. There is one color for regions with an even number of active neurons and another for regions with an odd number of active neurons. Because $f(x;\theta)$ lacks redundant neurons, there does not exist a boundary between activation regions where two neurons activations change simultaneously. In this work, we solely utilize this proposition for visualization purposes, as shown in Fig. 9. Nonetheless, we believe it may be of independent interest as it strengthens the connection between the surface of piecewise linear networks and the mathematics of paper folding, a connection previously alluded to in the literature [82].
@@ C.2 Dynamics of a two-layer piecewise linear network
We consider the dynamics of a two-layer piecewise linear network without biases, $f(x;\theta)=a^{\intercal}\sigma(Wx)$, where $W\in\mathbb{R}^{h\times d}$ and $a\in\mathbb{R}^{h}$. The activation function is $\sigma(z)=\max(z,\gamma z)$ for $\gamma\in[0,1)$, which includes ReLU $\gamma=0$ and Leaky ReLU $\gamma\in(0,1)$. We permit $h>d$, which in the limit as $h\to\infty$, ensures the network possesses the functional expressivity to represent any continuous nonlinear function from $\mathbb{R}^{d}$ to $\mathbb{R}$ passing through the origin. Following a similar strategy used in Section 4, we consider the contribution to the input-output map from a single hidden neuron $k\in[h]$ with parameters $w_{k}\in\mathbb{R}^{d}$ and $a_{k}\in\mathbb{R}$. As in the linear setting, each hidden neuron is associated with a conserved quantity, $\delta_{k}=\eta_{w}a_{k}^{2}-\eta_{a}\|w_{k}\|^{2}$. Unlike in the linear setting, this neuron’s contribution to the output $f(x_{i};\theta)$ is regulated by whether the input $x_{i}$ is in the neuron’s active halfspace, $\{x\in\mathbb{R}^{d}:w_{k}^{\intercal}x>0\}$. Let $C\in\mathbb{R}^{h\times n}$ be the matrix with elements $c_{ki}=\sigma^{\prime}(w_{k}^{\intercal}x_{i})$, which determines the activation of the $k^{\mathrm{th}}$ neuron for the $i^{\mathrm{th}}$ training data point. The subgradient $\sigma^{\prime}(z)=1$ if $z>0$, $\sigma^{\prime}(z)\in[\gamma,1]$ if $z=0$, and $\sigma^{\prime}(z)=\gamma$ if $z<0$. These activation functions exhibit positive homogeneity, implying $\sigma(z)=\sigma^{\prime}(z)z$. Thus, we can express $\sigma(w_{k}^{\intercal}x_{i})=c_{ki}w_{k}^{\intercal}x_{i}$, allowing us to express the gradient flow dynamics for $w_{k}$ and $a_{k}$ as
$\dot{a}_{k}=-\eta_{a}w_{k}^{\intercal}\left(\sum_{i=1}^{n}c_{ki}x_{i}\rho_{i}\right),\qquad\dot{w}_{k}=-\eta_{w}a_{k}\left(\sum_{i=1}^{n}c_{ki}x_{i}\rho_{i}\right),$ (128)
@@
where $\rho_{i}=f(x_{i};\theta)-y_{i}$ is the residual associated with the $i^{\mathrm{th}}$ training data point. If we let $\beta_{k}=a_{k}w_{k}$, which determines the contribution of each hidden neuron to the output $f(x_{i};\theta)$, then its not hard to see that the gradient flow dynamics of $\beta_{k}$ are
$\dot{\beta}_{k}=-\underbrace{\left(\eta_{w}a_{k}^{2}\mathbf{I}_{d}+\eta_{a}w_{k}w_{k}^{\intercal}\right)}_{M_{k}}\underbrace{\left(\textstyle\sum_{i=1}^{n}c_{ki}x_{i}\rho_{i}\right)}_{\xi_{k}}.$ (129)
@@
As in the linear setting, the matrix $M_{k}\in\mathbb{R}^{d\times d}$ appears as a preconditioning matrix on the dynamics Using the exact same derivation presented in Section A.3, whenever $a_{k}^{2}\neq 0$, we can express $M_{k}$ entirely in terms of $\beta_{k}$ and $\delta_{k}$,
$M_{k}=\frac{\sqrt{\delta_{k}^{2}+4\eta_{a}\eta_{w}\|\beta_{k}\|^{2}}+\delta_{k}}{2}\mathbf{I}_{d}+\frac{\sqrt{\delta_{k}^{2}+4\eta_{a}\eta_{w}\|\beta_{k}\|^{2}}-\delta_{k}}{2}\frac{\beta_{k}\beta_{k}^{\intercal}}{\|\beta_{k}\|^{2}}.$ (130)
@@
However, unlike in the linear setting, the vector $\xi_{k}\in\mathbb{R}^{d}$ driving the dynamics is not shared for all neurons because of its dependence on $c_{ki}$. Additionally, the NTK matrix in this setting depends on $M_{k}$ and $C$, with elements $K_{ij}=\sum_{k=1}^{h}c_{ki}x_{i}^{\intercal}\left(\eta_{w}a_{k}^{2}\mathbf{I}_{d}+\eta_{a}w_{k}w_{k}^{\intercal}\right)x_{j}c_{kj}$. Thus, in order to assess the temporal dynamics of the NTK matrix, we must understand the dynamics of $M_{k}$ and $C$. We consider a signed spherical coordinate transformation separating the dynamics of $\beta_{k}$ into its directional $\hat{\beta}_{k}=\mathrm{sgn}(a_{k})\tfrac{\beta_{k}}{\|\beta_{k}\|}$ and radial $\mu_{k}=\mathrm{sgn}(a_{k})\|\beta_{k}\|$ components, such that $\beta_{k}=\mu_{k}\hat{\beta}_{k}$. Here, $\hat{\beta}_{k}$ determines the orientation and direction of the halfspace where the $k^{\mathrm{th}}$ neuron is active, while $\mu_{k}$ determines the slope of the linear region in this halfspace. These coordinates evolve according to,
@@
$\dot{\mu}_{k}=-\sqrt{\delta_{k}^{2}+4\eta_{a}\eta_{w}\mu_{k}^{2}}\hat{\beta}_{k}^{\intercal}\xi_{k},\qquad\dot{\hat{\beta}}_{k}=-\frac{\sqrt{\delta_{k}^{2}+4\eta_{a}\eta_{w}\mu_{k}^{2}}+\delta_{k}}{2\mu_{k}}\left(\mathbf{I}_{d}-\hat{\beta}_{k}\hat{\beta}_{k}^{\intercal}\right)\xi_{k}.$ (131)
@@
These equations can be derived directly from Eq. 128 through chain rule similar to Section A.2.1. In fact its worth noting that the this change of coordinates is similar to the change of coordinates used in the single-neuron analysis. Expressed in terms of the parameters, $\hat{\beta}_{k}=\frac{w_{k}}{\|w_{k}\|}$ and $\mu_{k}=a_{k}\|w_{k}\|$.
@@ D.1 Figure 1: Teacher-Student with Two-layer ReLU Networks
Here the setting is: $d=100$, $m=50$, $k=3$, and $n=1000$, as in Fig. 1c of [8]. Training is performed with learning rate of 5e-3$/\tau^{2}$. Test error is computed as mean square error over a held-out set of 10,000 datapoints. We sweep over $\tau$ over a logarithmic scale in the range $[0.1,2]$ and $\delta$ over a linear scale in the range $[-1,1]$. We average over 16 random seeds, where the seed controls the sampling of: the teacher weights $\theta^{\mathrm{teacher}}$, the base initialization of $\theta^{\mathrm{student}}$, and the training data $\{x^{(i)}\}_{i=1}^{n}$. In this way, each random seed is used for a sweep over all combinations of $\tau$ and $\delta$ in the sweep; we simply apply the scaling described above to get to each point on the $(\tau,\delta)$ grid. The kernel distance computed is as defined in [27], where here we compute it at time $t$ relative to the kernel at initialization, i.e. $S(t)=1-\langle K_{0},K_{t}\rangle/\left(\|K_{0}\|_{F}\|K_{t}\|_{F}\right)$. In Fig. 10, we additionally plot Hamming and parameter distances relative to initialization, as well as training loss, while training for ten times longer than in Fig. 1 (b).
Notebooks generating all two-layer experiment figures are provided here.
@@ D.2 Figures 2, 3, 4: Single-Neuron Linear Network
Figures 2, 3, and 4 were generated by simulating gradient flow using scipy.integrate.solve_ivp function with the RK45 method for solving the ODEs, with a relative tolerance of $1\times 10^{-6}$ and time span of $(0,20)$. In the experiments with full-rank data, we used $X^{\intercal}X=\mathbf{I}_{2}$, $\beta_{*}=\left[\begin{smallmatrix}0\\ 1\end{smallmatrix}\right]$, and $\beta_{0}=\left[\begin{smallmatrix}-1\\ 0\end{smallmatrix}\right]$. For the experiment with low-rank data, we used $X^{\intercal}X=\left[\begin{smallmatrix}0.25&0.5\\ 0.5&1\end{smallmatrix}\right]$, $\beta_{*}=\left[\begin{smallmatrix}0.44\\ 0.88\end{smallmatrix}\right]$, and $\beta_{0}=\left[\begin{smallmatrix}0.4\\ 0.05\end{smallmatrix}\right]$. See the discussion in Section A.2 for details on how we determined our theoretical predictions. A notebook generating all the figures is provided here.
@@ D.3 Figure 5:
We are training a small ResNet based on the CIFAR10 script provided in the DAWN benchmark (code available here). The only modifications to the provided code base are we increase the convolution kernel size from $3\times 3$ to $15\times 15$, to better observe the learned spatial patterns, and we set the weight decay parameter to 00 to avoid confounding variables. Moreover, we are dividing the convolutional filters weights by a parameter $\alpha$ (after standard initialization) which controls the balancedness of the network. To quantify the smoothness of the filters, we compute the normalized Laplacian of each filter $w_{ij}\in\mathbb{R}^{15\times 15}$, over input $i=(1,2,3)$ and output $j=(1,...,64)$ channels
@@
We are training a one layer transformer model on the modular arithmetic task in Power et al. [68]. Our experimental code is based on an existing Pytorch implementation (code available here). The only modifications to the provided code base is that we use a single transformer layer (instead of the default 2-layer model). Prior analysis in Nanda et al. [72] has shown that this model can learn a minimal (attention-based) circuit that solves the task.